# Surgical SAM 2: Real-time Segment Anything in Surgical Video by Efficient Frame Pruning

### v1

@@ Abstract
Surgical video segmentation is a critical task in computer-assisted surgery and is vital for enhancing surgical quality and patient outcomes. Recently, the Segment Anything Model 2 (SAM2) framework has shown superior advancements in image and video segmentation. However, SAM2 struggles with efficiency due to the high computational demands of processing high-resolution images and complex and long-range temporal dynamics in surgical videos. To address these challenges, we introduce Surgical SAM 2 (SurgSAM-2), an advanced model to utilize SAM2 with an Efficient Frame Pruning (EFP) mechanism, to facilitate real-time surgical video segmentation. The EFP mechanism dynamically manages the memory bank by selectively retaining only the most informative frames, reducing memory usage and computational cost while maintaining high segmentation accuracy. Our extensive experiments demonstrate that SurgSAM-2 significantly improves both efficiency and segmentation accuracy compared to the vanilla SAM2. Remarkably, SurgSAM-2 achieves a 3$\times$ FPS compared with SAM2, while also delivering state-of-the-art performance after fine-tuning with lower-resolution data. These advancements establish SurgSAM-2 as a leading model for surgical video analysis, making real-time surgical video segmentation in resource-constrained environments a feasible reality.
Keywords:
Surgical Data Science Large Model Resource Efficient Learning Real-time Surgical Video Segmentation.
* These authors contribute equally to this work.Code will be publicly available at https://github.com/jinlab-imvr/Surgical-SAM-2.
@@ 1 Introduction
Surgical video scene segmentation is a critical task in computer-assisted surgery, where the precise identification and delineation of surgical instruments and tissues within video sequences are essential. This capability underpins various applications, such as instrument tracking and pose estimation, intraoperative guidance, and postoperative analysis [3, 4], ultimately enhancing surgical precision, reducing operative times, and improving patient outcomes. Achieving high segmentation accuracy with low computational cost is vital in real-world deployment considering limited resources in clinical centers. Real-time prediction also plays a core role in practical applications, to enable the method to provide timely decision-making support and navigation, generate real-time warning of potential deviations and anomalies, and facilitate remote surgical proctoring and team communication [1, 2]. However, accurate and efficient surgical video segmentation is challenging, due to the highly complicated surgical scene with lighting reflection and occlusion from blood and smoke, limited inter-class variance with highly similar appearance (e.g., different instruments), long-range duration with dynamic contexts, irregular artifacts, and noises.
@@
In the context of real-time surgical video segmentation, efficiency is not merely desirable but a critical necessity. Current approaches, including SAM2, have not adequately addressed this need, particularly in resource-constrained surgical environments where computational resources are limited. To address these challenges, we introduce Surgical SAM 2 (SurgSAM-2), one of the first works to specifically optimize SAM2 for surgical video segmentation, improving efficiency without compromising accuracy.
SurgSAM-2 is optimized for real-time application by integrating a dynamic memory bank management mechanism that reduces redundancy and computational load. This mechanism intelligently controls the retention of video frames by selectively pruning those that contribute less to ongoing analysis, thereby minimizing the impact of redundant frames that might otherwise dilute attention on critical objects. Unlike the original SAM2, which retains frames based on their order of arrival, SurgSAM-2 employs a dynamic scoring mechanism based on cosine similarity to assess the importance of each frame. By strategically reducing the memory bank size, SurgSAM-2 achieves state-of-the-art results, significantly increasing prediction speed in frames per second (FPS) and reducing memory requirements. This optimization ensures that SurgSAM-2 provides reliable and efficient performance in complex and dynamic scenario of surgical video analysis, making it a leading solution for real-time surgical video segmentation.
We conducted extensive experiments on instrument segmentation of surgical video, with two widely recognized benchmarks: the EndoVis17 and EndoVis18 datasets. The results demonstrate that SurgSAM-2 maintains high performance despite resource constraints, making it an ideal solution for real-time surgical video segmentation. This efficiency not only enables more accurate and timely decision-making during surgeries but also has the potential to reduce operating times and improve patient outcomes by ensuring that critical information is processed swiftly and accurately. Moreover, while our primary focus is on surgical video segmentation, the principles behind SurgSAM-2’s design have the potential to be adapted to broader video analysis tasks.
@@
    We introduce SurgSAM-2, a pioneering SAM2-based work for surgical video segmentation, designed to meet the specific needs of the surgical domain.
@@
    We extensively evaluate SurgSAM-2 on the EndoVis17 and EndoVis18 datasets, demonstrating its superior performance in both accuracy and efficiency of surgical instrument segmentation.
@@ 2.1 Surgical Instrument Segmentation
The field of surgical instrument segmentation has evolved significantly with deep learning, particularly through fully convolutional networks (FCNs) and encoder-decoder architectures like U-Net, which laid the foundation for the domain [15]. However, these early methods often faced challenges in dynamic surgical environments, struggling with spatial inconsistencies and the complex interactions between instruments and surrounding anatomy [16, 17].
To address these challenges, recent advancements have focused on transformer-based models and attention mechanisms, such as Swin Transformers and multi-scale attention U-Nets, which offer better robustness and adaptability in handling the complex visual features of surgical instruments [18, 19]. Following the introduction of the Segment Anything Model (SAM)[5] and SAM2[13], there has been a shift towards models tailored specifically for the medical image segmentation [39] and surgical video segmentation, such as SurgicalSAM [14]. Despite these advances, challenges remain, particularly in achieving efficient processing within resource constraints typical of surgical settings. This gap in efficiency is the core motivation behind SurgSAM-2, which seeks to optimize performance for real-time surgical applications [20, 21].
@@ 2.2 Segment Anything Model 2
SAM2 builds on Vision Transformers (ViTs) with enhanced multi-scale feature extraction, making it a powerful tool in image and video segmentation [13]. Following targeted modifications, SAM2 also shows substantial effectiveness in the 2D and 3D medical image segmentation [38]. However, its use in surgical video segmentation faces significant challenges due to the computational intensity of ViTs, which require substantial resources, limiting their practicality in real-time, resource-constrained environments [27]. SAM2’s reliance on a first-come-first-serve memory mechanism exacerbates inefficiencies, as it retains potentially redundant frames, further slowing down processing. The need for optimized models that reduce computational overhead while maintaining strong segmentation performance is critical, paving the way for more efficient solutions like SurgSAM-2.
@@ 2.3 Memory Bank Restriction
Efficient memory management is crucial for real-time applications, especially in the context of surgical video segmentation where computational resources are limited. Strategies like XMem [22] and RMem [23] have explored ways to prioritize and retain only the most relevant frames during video analysis. Building on these approaches, SurgSAM-2 introduces an efficient frame pruning mechanism, which uses a cosine similarity-based scoring system to retain only the most informative frames, reducing memory usage and increasing processing speed. This approach directly addresses inefficiencies in SAM2 memory management, making SurgSAM-2 better suited to the fast-paced and resource-constrained demands of real-time surgical video analysis.
@@ 3 Methods
SurgSAM-2 is an advanced model designed specifically for the complex and resource-constrained environments of surgical video segmentation. Building upon the SAM2 [13], SurgSAM-2 incorporates a dynamic memory bank management mechanism to optimize the retention and use of video frames during segmentation tasks. This innovation not only reduces the computational load but also enhances segmentation accuracy by selectively retaining the most relevant and useful information. The memory bank consists of the current frame and dynamically selected preceding frames, which are critical for maintaining temporal context. By integrating these advancements, SurgSAM-2 addresses the unique challenges posed by real-time surgical video analysis, offering a robust solution that balances efficiency and performance.
Figure 1: Architecture of the proposed model SurgSAM-2.
3.1 SurgSAM-2 Architecture
SurgSAM-2 builds upon the SAM2 [13], leveraging its robust ViT architecture specifically adapted for surgical video segmentation. The backbone of SAM2 is retained in SurgSAM-2 but with significant optimizations to meet the unique challenges posed by surgical environments.
At its core, the image encoder of SurgSAM-2 processes incoming video frames into detailed embeddings, capturing both local and global features necessary for accurate segmentation. While this architecture is consistent with the vanilla SAM2, SurgSAM-2 enhances it by integrating a dynamic memory management system that selectively prunes less relevant frames, ensuring that only the most critical data is retained for further analysis.
These refinements allow SurgSAM-2 to maintain the high performance associated with SAM2 while significantly improving efficiency, making it better suited for real-time applications in resource-constrained surgical settings.
@@ 3.2 Efficient Frame Pruning
A key innovation in SurgSAM-2 is the implementation of an efficient frame pruning mechanism, designed to intelligently manage which video frames are retained for further processing. This mechanism dynamically evaluates the relevance of each incoming frame before it is added to the memory bank, ensuring that only the most informative frames are preserved while discarding those that contribute minimally to the segmentation task.
@@
Considering that the vanilla SAM2 model uses a memory bank size of six past frames and the first reference frame for spatiotemporal modeling, our configuration is set as $n=5$ and $m=2$. We first compute the cosine similarity between frame $f_{t}$ and the past $n$ frames. Subsequently, we prune the two most similar frames $m=2$, leaving four frames in the dynamic memory bank, which aligns with our configuration for the EFP mechanism in SurgSAM-2.
By implementing this selective EFP strategy, SurgSAM-2 effectively reduces memory usage and computational load, allowing the model to process video frames more efficiently. In surgical videos, where scenes often exhibit high visual similarity and repeated semantic content across frames, this approach is particularly effective in eliminating redundancy while maintaining efficiency. Despite the reduction in stored frames, the model maintains high segmentation accuracy by focusing computational resources on the most critical data, making it well-suited for real-time surgical video analysis in resource-constrained environments.
@@ 4.1 Dataset
We extensively evaluate the proposed SurgSAM-2 model on two widely recognized and publicly available datasets: the 2017 MICCAI EndoVis Instrument Challenge (EndoVis17) [24] and the 2018 MICCAI EndoVis Scene Segmentation Challenge (EndoVis18) [25]. EndoVis18 presents more complicated surgical scenes, therefore proving more challenging compared to EndoVis17. The EndoVis17 dataset comprises eight training videos each containing 225 frames, eight testing videos collected followed by training videos, and another two hold-out testing videos (sequence 9, 10), comprising 1200 frames in total. The EndoVis18 dataset consists of 15 videos with each consisting of 149 frames. For Endovis17, we use the hold-out test set for evaluation. For EndoVis18, we split the sequences of 2, 5, 9, and 15 for testing following the standard procedure in ISINet [26].
@@ 4.2 Evaluation Metrics
To comprehensively evaluate the performance of SurgSAM-2, we adopt numerous and widely-used evaluation metrics in video object segmentation (VOS) that assess both the accuracy and computational efficiency of segmentation. These metrics are chosen to provide a holistic view of the model’s capabilities in the context of surgical video segmentation. For these metrics, we follow the evaluation protocol in the video object segmentation benchmark, specifically excluding the first and last frames from the assessment. We also utilize the official evaluation protocol in EndoVis Challenge [24, 25] for method validation.
@@
$J=\frac{\text{Area of Overlap}}{\text{Area of Union}}$
@@
$F=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
Dice=2×|PredictionGround Truth||Prediction|+|Ground $\text{Dice}=\frac{2\times|\text{Prediction}\cap\text{Ground Truth}|}{|\text{Prediction}|+|\text{Ground Truth}|}$
@@
Frames Per Second (FPS): To evaluate the real-time performance of SurgSAM-2, we measure the FPS during inference. This metric is crucial for applications in surgical environments where timely processing of video frames is essential for effective decision-making.
Memory Usage: We also assess the memory efficiency of SurgSAM-2 by calculating the model’s memory footprint during inference. Given the resource-constrained nature of many surgical settings, reducing memory usage is a key objective of our approach. By optimizing the memory bank size and employing selective frame retention, SurgSAM-2 achieves a balance between high segmentation accuracy and efficient memory usage.
These metrics together provide a comprehensive evaluation of SurgSAM-2, highlighting its strengths in segmentation accuracy, boundary precision, computational speed, and memory efficiency. By balancing these aspects, SurgSAM-2 is well-positioned to meet the demands of real-time surgical video analysis.
@@ 4.3.1 Evaluation on Model Efficiency
We conducted a thorough evaluation of SurgSAM-2’s performance in terms of FPS and memory usage, comparing it with the vanilla SAM2 across various configurations. The results, as illustrated in Table 1,  2 and 3, clearly demonstrate that by implementing a cosine similarity-based efficient frame pruning mechanism and reducing the memory bank size, both FPS and memory efficiency are significantly improved.
@@
In evaluating the efficiency of SurgSAM-2 compared to the vanilla SAM2, we observed consistent improvements across different prompt settings on both the EndoVis17 [24] and EndoVis18 [25] datasets. On average, SurgSAM-2 demonstrated a 13.8% increase in FPS and an 8.5% reduction in memory usage across the various prompt settings (Full Mask, One Point, and Five Points). These results underscore the effectiveness of our cosine similarity-based frame pruning mechanism in enhancing computational efficiency, particularly in resource-constrained environments.

### v2

@@ Abstract
Surgical video segmentation is a critical task in computer-assisted surgery and is vital for enhancing surgical quality and patient outcomes. Recently, the Segment Anything Model 2 (SAM2) framework has shown superior advancements in image and video segmentation. However, SAM2 struggles with efficiency due to the high computational demands of processing high-resolution images and complex and long-range temporal dynamics in surgical videos. To address these challenges, we introduce Surgical SAM 2 (SurgSAM2), an advanced model to utilize SAM2 with an Efficient Frame Pruning (EFP) mechanism, to facilitate real-time surgical video segmentation. The EFP mechanism dynamically manages the memory bank by selectively retaining only the most informative frames, reducing memory usage and computational cost while maintaining high segmentation accuracy. Our extensive experiments demonstrate that SurgSAM2 significantly improves both efficiency and segmentation accuracy compared to the vanilla SAM2. Remarkably, SurgSAM2 achieves a 3$\times$ FPS compared with SAM2, while also delivering state-of-the-art performance after fine-tuning with lower-resolution data. These advancements establish SurgSAM2 as a leading model for surgical video analysis, making real-time surgical video segmentation in resource-constrained environments a reality. Our source code is available at https://github.com/jinlab-imvr/Surgical-SAM-2.
@@ 1 Introduction
Surgical video scene segmentation is a critical task in computer-assisted surgery, where the precise identification and delineation of surgical instruments and tissues within video sequences are essential. This capability underpins various applications, such as instrument tracking and pose estimation, intraoperative guidance, and postoperative analysis [22, 9], ultimately enhancing surgical precision, reducing operative times, and improving patient outcomes. Achieving high segmentation accuracy with low computational cost is vital in real-world deployment considering limited resources in clinical centers. Real-time prediction also plays a core role in practical applications, to enable the method to provide timely decision-making support and navigation, generate real-time warning of potential deviations and anomalies, and facilitate remote surgical proctoring and team communication [10]. However, accurate and efficient surgical video segmentation is challenging, due to the highly complicated surgical scene with lighting reflection and occlusion from blood and smoke, limited inter-class variance with highly similar appearance (e.g., different instruments), long-range duration with dynamic contexts, irregular artifacts, and noises.
@@
In the context of real-time surgical video segmentation, efficiency is not merely desirable but a critical necessity. Current approaches, including SAM2, have not adequately addressed this need, particularly in resource-constrained surgical environments where computational resources are limited. To address these challenges, we introduce Surgical SAM 2 (SurgSAM2), one of the first works to specifically optimize SAM2 for surgical video segmentation, improving efficiency without compromising accuracy.
SurgSAM2 is optimized for real-time application by integrating a dynamic memory bank management mechanism that reduces redundancy and computational load. This mechanism intelligently controls the retention of video frames by selectively pruning those that contribute less to ongoing analysis, thereby minimizing the impact of redundant frames that might otherwise dilute attention on critical objects. Unlike the original SAM2, which retains frames based on their order of arrival, SurgSAM2 employs a dynamic scoring mechanism based on cosine similarity to assess the importance of each frame. By strategically reducing the memory bank size, SurgSAM2 achieves state-of-the-art results, significantly increasing prediction speed in frames per second (FPS) and reducing memory requirements. This optimization ensures that SurgSAM2 provides reliable and efficient performance in complex and dynamic scenario of surgical video analysis, making it a leading solution for real-time surgical video segmentation.
We conducted extensive experiments on instrument segmentation of surgical video, with two widely recognized benchmarks: the EndoVis17 and EndoVis18 datasets. The results demonstrate that SurgSAM2 maintains high performance despite resource constraints, making it an ideal solution for real-time surgical video segmentation. This efficiency not only enables more accurate and timely decision-making during surgeries but also has the potential to reduce operating times and improve patient outcomes by ensuring that critical information is processed swiftly and accurately. Moreover, while our primary focus is on surgical video segmentation, the principles behind SurgSAM2’s design have the potential to be adapted to broader video analysis tasks.
@@
    We introduce SurgSAM2, a pioneering SAM2-based work for surgical video segmentation, designed to meet the specific needs of the surgical domain.
@@
    We extensively evaluate SurgSAM2 on the EndoVis17 and EndoVis18, demonstrating its superior performance in both accuracy and efficiency of surgical instrument segmentation.
@@ 2.1 Surgical Instrument Segmentation
The field of surgical instrument segmentation has evolved significantly with deep learning, particularly through fully convolutional networks (FCNs) and encoder-decoder architectures like U-Net, which laid the foundation for the domain [23]. However, these early methods often faced challenges in dynamic surgical environments, struggling with spatial inconsistencies and the complex interactions between instruments and surrounding anatomy [18, 12, 19].
To address these challenges, recent advancements have focused on transformer-based models and attention mechanisms, such as Swin Transformers and multi-scale attention U-Nets, which offer better robustness and adaptability in handling the complex visual features of surgical instruments [15, 5]. Following the introduction of the Segment Anything Model (SAM)[14] and SAM2[21], there has been a shift towards models tailored specifically for the medical image segmentation [24] and surgical video segmentation, such as SurgicalSAM [27]. Despite these advances, challenges remain, particularly in achieving efficient processing within resource constraints typical of surgical settings. This gap in efficiency is the core motivation behind SurgSAM2, which seeks to optimize performance for real-time surgical applications [16, 17].
@@ 2.2 Segment Anything Model 2
SAM2 builds on Vision Transformers (ViTs) with enhanced multi-scale feature extraction, making it a powerful tool in image and video segmentation [21]. Following targeted modifications, SAM2 also shows substantial effectiveness in the 2D and 3D medical image segmentation [31]. However, its use in surgical video segmentation faces significant challenges due to the computational intensity of ViTs, which require substantial resources, limiting their practicality in real-time, resource-constrained environments [29]. SAM2’s reliance on a first-come-first-serve memory mechanism exacerbates inefficiencies, as it retains potentially redundant frames, further slowing down processing. The need for optimized models that reduce computational overhead while maintaining strong segmentation performance is critical, paving the way for more efficient solutions like SurgSAM2.
@@ 2.3 Memory Bank Restriction
Efficient memory management is crucial for real-time applications, especially in the context of surgical video segmentation where computational resources are limited. Strategies like XMem [8] and RMem [30] have explored ways to prioritize and retain only the most relevant frames during video analysis. Building on these approaches, SurgSAM2 introduces an efficient frame pruning mechanism, which uses a cosine similarity-based scoring system to retain only the most informative frames, reducing memory usage and increasing processing speed. This approach directly addresses inefficiencies in SAM2 memory management, making SurgSAM2 better suited to the fast-paced and resource-constrained demands of real-time surgical video analysis.
@@ 3 Methods
SurgSAM2 is an advanced model designed specifically for the complex and resource-constrained environments of surgical video segmentation. Building upon the SAM2 [21], SurgSAM2 incorporates a dynamic memory bank management mechanism to optimize the retention and use of video frames during segmentation tasks. This innovation not only reduces the computational load but also enhances segmentation accuracy by selectively retaining the most relevant and useful information. The memory bank consists of the current frame and dynamically selected preceding frames, which are critical for maintaining temporal context. By integrating these advancements, SurgSAM2 addresses the unique challenges posed by real-time surgical video analysis, offering a robust solution that balances efficiency and performance.
Figure 1: Architecture of the proposed model SurgSAM2.
3.1 SurgSAM2 Architecture
SurgSAM2 builds upon the SAM2 [21], leveraging its robust ViT architecture specifically adapted for surgical video segmentation. The backbone of SAM2 is retained in SurgSAM2 but with significant optimizations to meet the unique challenges posed by surgical environments.
At its core, the image encoder of SurgSAM2 processes incoming video frames into detailed embeddings, capturing both local and global features necessary for accurate segmentation. While this architecture is consistent with the vanilla SAM2, SurgSAM2 enhances it by integrating a dynamic memory management system that selectively prunes less relevant frames, ensuring that only the most critical data is retained for further analysis.
These refinements allow SurgSAM2 to maintain the high performance associated with SAM2 while significantly improving efficiency, making it better suited for real-time applications in resource-constrained surgical settings.
@@ 3.2 Efficient Frame Pruning
A key innovation in SurgSAM2 is the implementation of an efficient frame pruning mechanism, designed to intelligently manage which video frames are retained for further processing. This mechanism dynamically evaluates the relevance of each incoming frame before it is added to the memory bank, ensuring that only the most informative frames are preserved while discarding those that contribute minimally to the segmentation task.
@@
Considering that the vanilla SAM2 model uses a memory bank size of six past frames and the first reference frame for spatiotemporal modeling, our configuration is set as $n=5$ and $m=2$. We first compute the cosine similarity between frame $f_{t}$ and the past $n$ frames. Subsequently, we prune the two most similar frames $m=2$, leaving four frames in the dynamic memory bank, which aligns with our configuration for the EFP mechanism in SurgSAM2.
By implementing this selective EFP strategy, SurgSAM2 effectively reduces memory usage and computational load, allowing the model to process video frames more efficiently. In surgical videos, where scenes often exhibit high visual similarity and repeated semantic content across frames, this approach is particularly effective in eliminating redundancy while maintaining efficiency. Despite the reduction in stored frames, the model maintains high segmentation accuracy by focusing computational resources on the most critical data, making it well-suited for real-time surgical video analysis in resource-constrained environments.
@@ 4.1 Dataset
We extensively evaluate the proposed SurgSAM2 model on two widely recognized and publicly available datasets: the 2017 MICCAI EndoVis Instrument Challenge (EndoVis17) [2] and the 2018 MICCAI EndoVis Scene Segmentation Challenge (EndoVis18) [1]. EndoVis18 presents more complicated surgical scenes, therefore proving more challenging compared to EndoVis17. The EndoVis17 dataset comprises eight training videos each containing 225 frames, eight testing videos collected followed by training videos, and another two hold-out testing videos (sequence 9, 10), comprising 1200 frames in total. The EndoVis18 dataset consists of 15 videos with each consisting of 149 frames. For Endovis17, we use the hold-out test set for evaluation. For EndoVis18, we split the sequences of 2, 5, 9, and 15 for testing following the standard procedure in ISINet [10].
@@ 4.2 Evaluation Metrics
To comprehensively evaluate the performance of SurgSAM2, we adopt numerous and widely-used evaluation metrics in video object segmentation (VOS) that assess both the accuracy and computational efficiency of segmentation [20]. These metrics are chosen to provide a holistic view of the model’s capabilities in the context of surgical video segmentation. For these metrics, we follow the evaluation protocol in the video object segmentation benchmark, specifically excluding the first and last frames from the assessment. We also utilize the official evaluation protocol in EndoVis Challenge [2, 1] for method validation.
@@
Frames Per Second (FPS): To evaluate the real-time performance of SurgSAM2, we measure the FPS during inference. This metric is crucial for applications in surgical environments where timely processing of video frames is essential for effective decision-making.
Memory Usage: We also assess the memory efficiency of SurgSAM2 by calculating the model’s memory footprint during inference. Given the resource-constrained nature of many surgical settings, reducing memory usage is a key objective of our approach. By optimizing the memory bank size and employing selective frame retention, SurgSAM2 achieves a balance between high segmentation accuracy and efficient memory usage.
These metrics together provide a comprehensive evaluation of SurgSAM2, highlighting its strengths in segmentation accuracy, boundary precision, computational speed, and memory efficiency. By balancing these aspects, SurgSAM2 is well-positioned to meet the demands of real-time surgical video analysis.
@@ 4.3.1 Evaluation on Model Efficiency
We conducted a thorough evaluation of SurgSAM2’s performance in terms of FPS and memory usage, comparing it with the vanilla SAM2 across various configurations. The results, as illustrated in Table 1,  2 and 3, clearly demonstrate that by implementing a cosine similarity-based efficient frame pruning mechanism and reducing the memory bank size, both FPS and memory efficiency are significantly improved.
@@
In evaluating the efficiency of SurgSAM2 compared to the vanilla SAM2, we observed consistent improvements across different prompt settings on both the EndoVis17 [2] and EndoVis18 [1] datasets. On average, SurgSAM2 demonstrated a 13.8% increase in FPS and an 8.5% reduction in memory usage across the various prompt settings (Full Mask, One Point, and Five Points). These results underscore the effectiveness of our cosine similarity-based frame pruning mechanism in enhancing computational efficiency, particularly in resource-constrained environments.